Miranda wants to make 16L of a 50% saline solution by mixing together a 20% saline solution and a 60% saline solution. How much of each solution must she use?
Answer by lwsshak3(11628)

Miranda wants to make 16L of a 50% saline solution by mixing together a 20% saline solution and a 60% saline solution. How much of each solution must she use?
***
let x=amt of 20% solution to use
16-x=amt of 60% solution to use
..
sum of saline in indv. solutions=amt of saline in mixed solution
20%x+60%(16-x)=50%*16
.20x+9.6-.60x=8
.40x=1.6
x=4
16-x=12
amt of 20% solution to use=4L
amt of 60% solution to use=12L